# Lightweight Models for Emotional Analysis in Video

### v1

@@ 1 Introduction
The challenge includes six tasks: Valence-Arousal (VA) Estimation, Expression (EXPR) Classification, Action Unit(AU) Detection, Compound Expression (CE) Recognition, Emotional Mimicry Intensity (EMI) Estimation, Ambivalence/Hesitancy (AH) Recognition. These challenges leverage datasets such as Aff-Wild2, C-EXPR-DB, HUME-Vidmimic2, and BAH, providing a comprehensive benchmark for evaluating affective behavior analysis models.The challenges utilize the Aff-Wild2 [8, 11], C-EXPR-DB [7], HUME-Vidmimic2, and BAH datasets, providing a comprehensive evaluation framework for affective behavior analysis methodologies. In this paper, we solve the tasks such as Valence-Arousal (VA) Estimation, Action Unit(AU) Detection, Emotional Mimicry Intensity (EMI) Estimation, Ambivalence/Hesitancy (AH) Recognition. The VA task predicts valence and arousal in video sequences, while the AU task detects the presence of 12 action units (AUs) in each video frame. The EMI task estimates the intensity of six emotional dimensions (Admiration, Amusement, Determination, Empathic Pain, Excitement, and Joy). Lastly, the AH task identifies the presence or absence of ambivalence or hesitancy in each frame.
@@ 3.1 Dataset
This challenge uses a data set that includes 542 videos with annotations for 12 Action Units (AU), which represent facial muscle movements, including brow raisers, cheek raisers, lip tighteners, and jaw drops. The data set consists of 2,627,632 frames captured from 438 unique subjects. Annotations were developed through a semiautomatic methodology that integrates both manual and computational techniques. The data set has been divided into three subsets: a training set (295 videos), a validation set (105 videos), and a testing set (142 videos). Table 1 provides the distribution of the AU annotations of the dataset.
Table 1: Distribution of AU Annotations in Aff-Wild2
AU Action Total Number of
Activated AUs
AU 1 inner brow raiser 301,102
AU 2 outer brow raiser 139,936
AU 4 brow lowerer 386,689
AU 6 cheek raiser 619,775
AU 7 lid tightener 964,312
AU 10 upper lip raiser 854,519
AU 12 lip corner puller 602,835
AU 15 lip corner depressor 63,230
AU 23 lip tightener 78,649
AU 24 lip pressor 61,500
AU 25 lips part 1,596,055
AU 26 jaw drop 206,535
@@
Table 3: Action unit validation set.
Methods F1 score
Regnet-ViT [23] 0.5280
EmotiEffNet [15] 0.537
Ours with 3 level mixer 0.5369
Ours with 2 level mixer 0.5338
Ours - single mixer 0.5441

### v2

@@ 1 Introduction
The challenge includes six tasks: Valence-Arousal (VA) Estimation, Expression (EXPR) Classification, Action Unit(AU) Detection, Compound Expression (CE) Recognition, Emotional Mimicry Intensity (EMI) Estimation, Ambivalence/Hesitancy (AH) Recognition. The challenges utilize the Aff-Wild2 [35, 20, 19, 18, 17, 9, 8, 7, 12, 11, 16, 13, 10, 15, 14], C-EXPR-DB [9], HUME-Vidmimic2, and BAH datasets, providing a comprehensive evaluation framework for affective behavior analysis methodologies. In this paper, we solve the tasks such as Valence-Arousal (VA) Estimation, Action Unit(AU) Detection, Emotional Mimicry Intensity (EMI) Estimation, Ambivalence/Hesitancy (AH) Recognition. The VA task predicts valence and arousal in video sequences, while the AU task detects the presence of 12 action units (AUs) in each video frame. The EMI task estimates the intensity of six emotional dimensions (Admiration, Amusement, Determination, Empathic Pain, Excitement, and Joy). Lastly, the AH task identifies the presence or absence of ambivalence or hesitancy in each frame.
@@ 3.1 Dataset
This challenge uses a data set that includes 542 videos with annotations for 12 Action Units (AU), which represent facial muscle movements, including brow raisers, cheek raisers, lip tighteners, and jaw drops. The data set consists of 2,627,632 frames captured from 438 unique subjects. Annotations were developed through a semiautomatic methodology that integrates both manual and computational techniques. The data set has been divided into three subsets: a training set (295 videos), a validation set (105 videos), and a testing set (142 videos). LABEL:tab:au_dis provides the distribution of the AU annotations of the dataset.
@@
Table 2: Evaluation on Action Unit detection task with F1 score.
Methods Validation Test
Regnet-ViT [34] 0.5280 -
EmotiEffNet [25, 26] 0.537 0.488
AIWELL-UOC [2] 0.451 0.417
Ours with 3 level mixer 0.5369 -
Ours with 2 level mixer 0.5338 -
Ours - single mixer 0.5441 0.484